INFRARED SPECTROSCOPY OF PHENOL-TRIETHYLSILANE DIHYDROGEN-BONDED CLUSTER

\underline{HARUKI ISHIKAWA}, TAKAYUKI KAWASAKI, Department of Chemistry, School of Science, Kitasato University, Minami-ku, Sagamihara 252-0373, Japan.

Dihydrogen bond is a hydrogen bond between oppositely charged two hydrogen atoms, X-H⋯H-Y, where X = O, N and Y = B, metal atoms, for example. In 2005, Ishikawa and coworkers reported the observation of the dihydrogen-bond system involving Si-H group as the proton acceptor^{a}. They carried out infrared (IR) spectroscopy of phenol(PhOH)- Diethylmethylsilane(DEMS) clusters. All of the three isomers of PhOH-DEMS 1:1 clusters observed exhibit a small red-shift of ν̃_{OH} of the PhOH moiety in the cluster compared with that of bare PhOH. The largest shift is −29 cm^{−1}. The small red-shift is considered to be the result of the competition between the O-H⋯H-Si dihydrogen-bond and the dispersion interaction of alkyl group of DEMS with phenyl ring. It means that the strength of the O-H⋯H-Si dihydrogenbond is comparable to the dispersion force. In the present study, we have performed fluorescence excitation (FE) and IR spectroscopies of phenol-triethylsilane(TES) to widen the knowledge of the dihydrogen bond. Similar to the case of PhOH-DEMS system, the electronic origin bands of three PhOH-TES isomers appear in the vicinity of that of PhOH monomer in the FE spectrum. In the present study, we have found an origin band of another PhOH-TES isomer showing a red-shift of −120 cm^{−1}. The shift of ν̃_{OH} of this cluster is found to be −78 cm^{−1}. This value is much larger than those of the other PhOH-TES 1:1 clusters. It is expected that the spatial overlap of between the TES and the phenyl ring in this cluster is small so that the contribution of the O-H⋯H-Si dihydrogen-bond becomes larger than the other isomers. We have performed density-functional-theory (DFT) calculation of the PhOH-TES clusters using M05-2X functional. The result of the DFT calculation supported the cluster structure and the large red-shift of ν̃_{OH} of the newly found isomer of PhOH-TES.

aH. Ishikawa, A. Saito, M. Sugiyama, N. Mikami, J. Chem. Phys. 123, 224309 (2005).